OpenStudy (anonymous):

I need someone to explain to me step by step the following problem:

OpenStudy (anonymous):

Find the equation of the tangent line to the curve \[x^{2}y^{2}-1=3(x+y)\] at the points (1,-1)

OpenStudy (anonymous):

I have the problem worked out in front of me (prof gave us the answers) but, his work is disorganized and I don't understand what it is he's doing.

OpenStudy (anonymous):

I figure I have to do this implicitly.

OpenStudy (asnaseer):

yes - you are correct - use implicit differentiation to get an expression for slope of the tangent line at any point (x,y). then substitute x=1, y=-1 to find the value of the slope of the tangent line at (1, -1) then use the equation of a straight line: y = mx + c to find c (since you now know m) by using the fact that it passes through the point (1, -1)

OpenStudy (anonymous):

is the answer y=4-5x?

OpenStudy (anonymous):

that's not the answer.... the answer is: y=-1/5x+4/5

OpenStudy (anonymous):

its a plane....?

OpenStudy (anonymous):

okay and the way my prof did his differentiation was really weird. his first step looks like: x^{2}2y*y'+y^{2}(2x)=3(1+y')

OpenStudy (anonymous):

a plane? huh?

OpenStudy (anonymous):

that's not how i would start the implicit differentiation.

OpenStudy (anonymous):

\[x^{2}(2y*y')+y^{2}(2x)=3(1+y')\]

OpenStudy (asnaseer):

what your prof did is correct - he is using the product rule, see here: http://www.1728.org/chainrul.htm

OpenStudy (asnaseer):

how would you have done this?

OpenStudy (anonymous):

haha no idea... i suck at math : (

OpenStudy (anonymous):

so i use the product rule for x^2*y^2 ??

OpenStudy (asnaseer):

correct

OpenStudy (anonymous):

and for the 3*(x+y) ??

OpenStudy (asnaseer):

that is just straight differentiation since it can be written as: 3x + 3y

OpenStudy (anonymous):

so that will just equal 3y' ??

OpenStudy (anonymous):

or 3+3y'

OpenStudy (asnaseer):

so you can effectively just ignore the 3 outside the braces and just differentiate what is inside the braces

OpenStudy (anonymous):

so it's 3(1+y')

OpenStudy (asnaseer):

yes - correct

OpenStudy (anonymous):

okay so now i have the same first line as he does haha.

OpenStudy (anonymous):

i ignored the 3 outside the brackets because it's a constant right?

OpenStudy (asnaseer):

yes. thats good - you are making progress - maybe you are a lot better at maths than you think. :)

OpenStudy (anonymous):

haha maybe, maybe not...

OpenStudy (anonymous):

so now i can take the y^2(2x) over to the other side?

OpenStudy (anonymous):

and the right hand side over to the left?

OpenStudy (asnaseer):

yes - what you need to do is to get an expression for y'. so move all terms involving y' to the left and the rest to the right

OpenStudy (asnaseer):

first expand the expression on the right

OpenStudy (anonymous):

do i leave my 3 on the right hand side?

OpenStudy (asnaseer):

snap! yes :)

OpenStudy (anonymous):

okay so now i have: (x^2)(2y*y')-(1-y')=3(-2xy^2)

OpenStudy (anonymous):

that look right to you?

OpenStudy (asnaseer):

not quite :(

OpenStudy (anonymous):

oh snappp

OpenStudy (asnaseer):

you started with this:\[x^{2}(2y*y')+y^{2}(2x)=3(1+y')\]so first expand the right to get:\[x^{2}(2y*y')+y^{2}(2x)=3+3y'\]then move terms involving y' to the left and the rest to the right to get:\[x^{2}(2y*y')-3y'=3-y^{2}(2x)\]

OpenStudy (asnaseer):

then factorise y' on the left

OpenStudy (anonymous):

\[y'=\frac{ 3-2xy^{2} }{ 2x^{2}y-3 }\]

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

woooo. so now that part is done, right?

OpenStudy (asnaseer):

yup - the rest should be straight forward. I need to go for dinner now. I'll check back afterwards to see if you had any further problems.

OpenStudy (asnaseer):

the steps left are: then substitute x=1, y=-1 to find the value of the slope of the tangent line at (1, -1) then use the equation of a straight line: y = mx + c to find c (since you now know m) by using the fact that it passes through the point (1, -1)

OpenStudy (anonymous):

okay thanks. : )

OpenStudy (asnaseer):

yw :) and good luck!

OpenStudy (anonymous):

i got the rest of it now ; ) thanks a bunch!

OpenStudy (asnaseer):

thats great! and you are more than welcome! :)